A body covers 10 metre in second second and 25 metre in fifth second of its motion . If the motion is uniformly accelerated. Then how far it will go in 7th second.

Grade:12

1 Answers

Janhavi Srivastava
6 years ago
S nth= u+a/2(2n-1) 10=u+a/2(2*2-1)=10= u+3a/2..........(1)25=u+a/2(5*5-1)=25=u+9a/2...........(2)From (1) and (2),a=5m/s2And u=2.5m/sFor 7th sec,S7= 2.5+5/2*(7*2-1) =2.5+32.5 = 35mThank you